# Multi-Scale Attention and Gated Shifting for Fine-Grained Event Spotting in Videos

### v1

@@ 4 Empirical Study
Figure 5: Example frames from the proposed TTA dataset. The top row shows samples from World Para Table Tennis (WPF) matches, while the bottom row displays clips from the 2024 Paris Paralympic Games. The dataset captures diverse viewing angles, frequent occlusions, and high event density representative of real-world elite-level table tennis.
In this section, we first review all the datasets used to evaluate the proposed MSAGSM, with a detailed description of the newly introduced TTA dataset. We then outline the implementation details of our method, conduct an ablation study to analyze the contribution of each module component, and finally present a comparative evaluation against prior methods and current state-of-the-art approaches.
@@ 4.1 Datasets
We evaluate our method on four fine-grained sports video datasets with frame-level annotations: Tennis [40], Figure Skating [10], FineDiving [38], and our proposed Table Tennis Australia (TTA) dataset.
The Tennis dataset consists of 3,345 video clips from 28 matches, with frame rates of 25 or 30 FPS. It contains 33,791 frame-accurate annotations spanning six event classes. The Figure Skating dataset includes 11 broadcast videos (all 25 FPS), covering 371 short program performances from major international competitions between 2010 and 2019. It includes 3,674 annotated events across four classes. Following previous work [10, 35, 29], we adopt two evaluation splits: a competition split (FS-Comp) and a performance split (FS-Perf). The FineDiving dataset contains 3,000 diving clips with temporal segment annotations. These have been converted into frame-level event labels across four classes using the annotation protocol from [11].
We also introduce the Table Tennis Australia (TTA) dataset, the first precise event spotting benchmark for table tennis. It consists of 39 full-game videos recorded at 30 FPS, totaling 4,878 annotated events across 8 classes: serve, bounce, forehand, and backhand, each labeled separately for near and far sides of the table. Examples of the TTA dataset are presented in Figure 5. Events are annotated following labeling protocols from Table Tennis Australia. For example, a frame is labeled as an event when the ball contacts the paddle or undergoes significant motion immediately after. In cases of occlusion (e.g., when a player blocks the ball near the camera), temporal context from preceding and following frames is used to estimate the most accurate event frame. For more details, please refer to the supplementary materials.
@@
Table 1: Performance comparison (mAP %) under different temporal tolerances ($\delta=1$ and $\delta=2$). Results are reported on five fine-grained sports video datasets. The best result for each column is bolded, and our proposed method is italicized. $\star$ indicates the best result within each model-backbone group. All models are evaluated using a fixed clip length of 100 frames.
Model Backbone Temp Module TTA Tennis FS_Comp FS_Perf FineDiving Parameters
$\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2 $\delta$=1 $\delta$=2
E2E-Spot RegNetY-200 GSM 66.42 76.90 90.03 96.48 75.39 89.33 81.27 94.14 67.89 86.97 4.45M
E2E-Spot RegNetY-200 GSF 63.70 72.19 89.01 96.55 75.49 89.32 81.55 94.02 67.57 86.22 4.45M
E2E-Spot RegNetY-200 Ours 69.50 78.12 91.34 96.51 77.72 90.04 84.12 96.90 71.03 87.26 4.50M
T-DEED RegNetY-200 GSM 60.69 69.30 89.21 97.15 71.69 90.02 66.18 84.02 66.18 84.02 16.42M
T-DEED RegNetY-200 GSF 54.04 64.51 89.58 97.06 75.52 92.15 79.82 94.79 67.04 83.24 16.42M
T-DEED RegNetY-200 Ours 63.52 71.16 90.68 96.67 75.55 87.61 82.41 94.12 66.70 84.60 16.48M
UGLF RegNetY-200 GSM 63.52 71.16 89.73 97.16 77.25 90.44 79.24 93.08 66.88 84.94 4.45M
UGLF RegNetY-200 GSF 63.70 72.19 89.39 97.06 72.61 86.90 82.06 94.38 66.65 84.27 4.45M
UGLF RegNetY-200 Ours 64.82 74.84 90.98 97.10 78.86 91.13 82.64 96.12 67.21 85.34 4.50M
Unlike existing datasets, TTA videos are collected using handheld cameras by professional analysts from Paralympics Australia. The dataset includes elite-level competitions such as the Paralympics and WPF. This real-world, single-camera setting introduces challenges such as frequent occlusions, motion blur, and dynamic angles—conditions closer to practical applications—unlike the broadcast-quality videos used in most prior work. As shown in Figure 6, TTA exhibits the highest event density among all datasets, averaging 7.73 events per 100-frame window, compared to 3.81 in Tennis and significantly fewer in others. This emphasizes the need for models that can capture fine-grained temporal structure and distinguish between visually similar frames within narrow time intervals.
@@ 4.2 Evaluation
Since our proposed module, MSAGSM, is designed to be a lightweight and drop-in replacement for GSM, we follow the same training procedure as the E2E-Spot baseline [11], replacing GSM with MSAGSM in the backbone. The model performance is evaluated using the mean Average Precision (mAP), computed by averaging the Average Precision (AP) across all event classes. Following prior work, we report results under two temporal tolerance settings: $\delta=1$ and $\delta=2$, representing correct prediction within 1 frame distance and 2 frame distances.
@@ 4.3 Implementation Details
We follow the training protocol established in E2E-Spot. All models are trained using input clips of length $L=100$ frames and a batch size of 8. Each epoch comprises 5,000 clips randomly sampled from the training set. Training is conducted for 50 epochs using the AdamW optimizer [18] with a base learning rate of $1\times 10^{-3}$. We apply 3 warm-up epochs followed by a cosine decay schedule. To address class imbalance, positive classes in the cross-entropy loss are weighted by a factor of $w=5$. We evaluate our MSAGSM module using two backbone architectures—RegNetY-200 [21] and ResNet-18 [9]—both adopted from the original E2E framework. All experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.4 Comparison with State-of-the-Art
To assess the effectiveness and generalizability of MSAGSM, we compare it with three representative models in the PES literature: E2E-Spot, T-DEED, and UGLF. All three models follow a common architectural structure: a 2D backbone for feature extraction, followed by a temporal modeling module, and a decoder for frame-level prediction. E2E-Spot serves as a strong baseline due to its simplicity and broad adoption in the field, while T-DEED and UGLF build upon E2E by introducing additional modules or input features to model long-range dependencies or enhance feature discrimination (e.g., via SGP, GLIP). Each method adopts a different temporal modeling choice—E2E and UGLF both report GSM as their best-performing module, while T-DEED employs GSF. Additionally, all three models use RegNetY-200 as their 2D backbone of choice, based on reported performance in their original implementations. To ensure fairness, we reused the officially released codebases and trained each model on our hardware under consistent settings. For each, we report the best performance across multiple runs.
@@
Main Results. Table 1 presents results on five fine-grained sports video datasets under two temporal tolerance thresholds ($\delta=1$ and $\delta=2$). Across the board, our proposed MSAGSM module consistently outperforms existing temporal modeling approaches. For example, within the E2E architecture using RegNetY-200, MSAGSM improves upon GSM by +3.08 mAP on TTA at $\delta=1$, and by +1.15 on FineDiving at $\delta=1$. Similar performance gains are observed when MSAGSM replaces the original temporal module in more complex models such as T-DEED and UGLF.
@@
Out of 30 possible comparisons across model-backbone configurations, MSAGSM achieves the best result in 24 cases when compared directly against other temporal modules. Furthermore, it ranks first in 8 out of 10 columns when considering overall best scores, setting new state-of-the-art performance simply by replacing the temporal component.
@@
Backbone and Module Versatility. To evaluate the generality of our approach, we further experiment with alternative backbones—specifically ResNet-18—in E2E, T-DEED, and UGLF architectures. For simplicity, the main table includes RegNetY-200 results only, while additional results for E2E, T-DEED and UGLF with ResNet-18 are provided in the supplementary materials. We also evaluate alternative temporal modules not originally reported in the respective papers: GSF in E2E and UGLF, and GSM in T-DEED. Across all these configurations, MSAGSM consistently delivers strong performance without increasing model complexity, demonstrating its robustness and adaptability. For both E2E-Spot and UGLF, which are relatively simpler models, our MSAGSM module demonstrates strong improvements over their original temporal module designs. Specifically, E2E-Spot with MSAGSM outperforms all other temporal modules across all five datasets, while UGLF with MSAGSM achieves the best performance in 9 out of 10 groups. We observe that the improvement is less pronounced in the T-DEED model, which is reasonable given that T-DEED already incorporates a more complex temporal modeling design. In such cases, introducing MSAGSM may add redundant or conflicting temporal information, which could complicate effective feature retrieval.
Setting Value TTA FineDiving Params
mAP@1 mAP@2 mAP@1 mAP@2
Varying Number of Attention Heads (Dilations = [1,2,3])
Heads 1 60.99 75.38 66.36 83.47 4.50M
Heads 2 69.50 (+8.51) 78.12 (+2.74) 71.03 (+4.67) 87.26 (+3.79) 4.51M
Heads 3 65.87 (+4.88) 75.87 (+0.49) 67.93 (+1.57) 85.11 (+1.64) 4.51M
Heads 4 61.01 (+0.02) 72.98 (–2.40) 68.72 (+2.36) 85.16 (+1.69) 4.52M
Heads 8 64.19 (+3.20) 72.94 (–2.44) 70.06 (+3.70) 85.91 (+2.44) 4.55M
@@
Distance [1,2] 60.78 71.10 68.56 85.49 4.48M
Distance [1,2,3] 69.50 (+8.72) 78.12 (+7.02) 71.03 (+2.47) 87.26 (+1.77) 4.51M
Distance [1,…,4] 64.18 (+3.40) 71.59 (+0.49) 69.02 (+0.46) 85.17 (-0.32) 4.53M
Distance [1,…,5] 67.54 (+6.76) 67.41 (-3.69) 69.26 (+0.70) 85.26 (-0.23) 4.55M
Distance [1,…,6] 64.03 (+3.25) 75.38 (+4.28) 69.84 (+1.28) 85.34 (-0.15) 4.57M
Efficiency. MSAGSM maintains a lightweight footprint, using only 4.5M parameters in its default configuration. This matches or only slightly exceeds the size of existing temporal modules, while consistently delivering superior accuracy. Its compact design makes MSAGSM a practical drop-in replacement for both lightweight and large-scale models that rely on 2D pretrained CNN backbones.
Summary. These results demonstrate the effectiveness of MSAGSM as a robust and efficient temporal modeling module. Its consistent improvements across diverse models and baseline temporal modules, particularly on our proposed TTA benchmark (see Figure 6), underscore its suitability for capturing the high-frequency, fine-grained temporal dynamics inherent in sports video understanding.
@@ 4.5 Ablation Study
We conduct ablation studies on the TTA and FineDiving datasets, which represent two contrasting types of sports—one with high event density and the other with sparse events. These experiments evaluate two key components of our proposed MSAGSM: the number of attention heads and the temporal window size (i.e., dilation range).
Number of Attention Heads. We fix the dilation setting to $[1,2,3]$ and vary the number of attention heads. As shown in Table 6, using two heads achieves the best trade-off across both datasets. On TTA, increasing from one to two heads substantially improves mAP by +8.51 and +2.74 at $\delta=1$ and $\delta=2$, respectively. However, increasing the number of heads beyond two leads to inconsistent or diminished performance, likely due to over-fragmentation of attention across feature subspaces. On FineDiving, which contains smoother and more globally distributed motion, two heads also yield the best result (71.03 mAP@1), with additional heads offering minimal or no further gains.
Temporal Window Size. We fix the number of heads at two and vary the dilation configurations to control the size of the temporal receptive field. The default setting $[1,2,3]$ performs best overall. Slight improvements are observed on TTA with longer ranges (e.g., $[1,2,3,4]$, +3.40 mAP@1), but extremely long windows such as $[1,\dots,6]$ result in diminishing or inconsistent gains. This suggests that overly large temporal windows may dilute fine-grained event cues. These findings highlight the importance of balancing broad context modeling with temporal precision.
In summary, MSAGSM achieves strong and stable performance with two attention heads and a temporal window of $[1,2,3]$, demonstrating its robustness across datasets with diverse temporal dynamics. Additional ablation results, including further window configurations, are provided in the supplementary materials.
5 Discussion
In this paper, we presented MSAGSM, a plug-and-play temporal module designed for integration with 2D backbone feature extractors in temporally stringent event spotting tasks. MSAGSM is a simple yet effective component that enhances models following the E2E-Spot framework [11], enabling improved performance—particularly on temporally precise spotting tasks in fine-grained sports videos. Our method consistently outperforms existing temporal modules such as GSM and GSF across various settings, including different backbones (e.g., RegNetY-200, ResNet-18) and model architectures (e.g., T-DEED, UGLF). Details and additional experiments, including results on ResNet-18 backbones, are provided in the supplementary materials.

### v2

@@ 4 Empirical Study
In this section, we describe the datasets (with details on the new TTA dataset), outline implementation settings, present ablation studies, and compare MSAGSM with prior methods and state-of-the-art baselines.
@@ 4.1 Datasets
We evaluate our method on four fine-grained sports video datasets with frame-level annotations: Tennis [53], Figure Skating [11], FineDiving [48], and our proposed TTA dataset.
Figure 5: Example frames from the proposed TTA dataset. The top row shows samples from World Para Future matches, while the bottom row displays clips from the 2024 Paris Paralympic Games. The dataset captures diverse viewing angles, frequent occlusions, and high event density, representative of real-world table tennis matches.
The Tennis dataset consists of 3,345 video clips from 28 matches, with frame rates of 25 or 30 FPS. It contains 33,791 frame-accurate annotations spanning six event classes. The Figure Skating dataset includes 11 broadcast videos (all 25 FPS), covering 371 short program performances from major international competitions between 2010 and 2019. It includes 3,674 annotated events across four classes. Following previous work [11, 44, 37], we adopt two evaluation splits: a competition split (FS-Comp) and a performance split (FS-Perf). The FineDiving dataset comprises 3,000 diving clips with temporal segment annotations, totaling 7,010 events. These have been converted into frame-level event labels across four classes using the annotation protocol from [12].
We also introduce the Table Tennis Australia (TTA) dataset details shown in Table 1, the first PES benchmark for table tennis. It consists of 39 full-game videos recorded at 30 FPS, totaling 4,878 annotated events across 8 classes: serve, bounce, forehand, and backhand, each labeled separately for near and far sides of the table. Examples of the TTA dataset are presented in Figure 5. Events are annotated following labeling protocols from Table Tennis Australia. For example, a frame is labeled as an event when the ball contacts the paddle or undergoes significant motion immediately after. In cases of occlusion (e.g., when a player blocks the ball near the camera), temporal context from preceding and following frames is used to estimate the most accurate event frame. For more details, please refer to the Supp. §2.
Table 1: Comparison of PES benchmark datasets. TTA (Bold) offers higher event density and realistic recording conditions.
Dataset Recording Events Classes Density
TTA (Ours) Handheld 4,878 8 7.73
Tennis Broadcast 33,791 6 3.81
Figure Skating Broadcast 3,674 4 1.58
FineDiving Broadcast 7010 4 1.49
@@
Table 2: Performance comparison (mAP %) at different temporal tolerances ($\delta=0,1,2$) across five fine-grained sports video datasets. The best result in each column is shown in bold, and our proposed method is italicized. A $\star$ marks the best result within each model–backbone group. All models are evaluated with a fixed clip length of 100 frames.
Model Temp TTA Tennis FS_Comp FS_Perf FineDiving P (M) / GFLOPs $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 $\delta$=0 $\delta$=1 $\delta$=2 ASTRA[43] - 28.12 54.16 60.95 49.98 88.67 97.49 35.87 73.45 84.71 39.35 78.10 90.05 27.28 65.01 82.74 4.49 / 42.20 ASFormer[50] - 21.68 40.69 61.08 42.57 89.34 93.39 32.60 50.29 55.97 16.34 36.46 58.41 15.86 66.82 87.11 3.21 / 41.94 E2E-Spot[12] GSM 34.36 66.42 76.90 51.42 90.03 96.48 36.44 75.39 89.33 39.32 81.27 94.14 27.31 67.89 86.97 4.45 / 42.18 E2E-Spot GSF 36.17 63.70 72.19 49.68 89.01 96.55 36.25 75.49 89.32 40.07 81.55 94.02 29.41 67.57 86.22 4.45 / 42.20 E2E-Spot Ours 36.43 69.50 78.12 55.51 91.34 96.51 38.34 77.72 90.04 46.54 84.12 96.90 29.83 71.03 87.26 4.51 / 45.06 T-DEED GSM 33.32 60.69 69.30 48.82 89.21 97.15 38.61 71.69 90.02 42.53 79.62 90.36 26.86 66.18 84.02 16.42 / 43.34 T-DEED[44] GSF 26.27 54.04 64.51 50.08 89.58 97.06 38.83 75.52 92.15 38.89 79.82 94.79 28.50 67.04 83.24 16.42 / 43.34 T-DEED Ours 28.06 60.77 72.50 52.76 90.68 96.67 39.06 75.55 87.61 44.66 82.41 94.12 26.88 66.70 84.60 16.48 / 46.20 UGLF[37] GSM 34.36 63.52 71.16 50.31 89.73 97.16 36.99 77.25 90.44 38.29 79.24 93.08 27.11 66.88 84.94 4.45 / 42.18 UGLF GSF 36.17 63.70 72.19 50.94 89.39 97.06 34.84 72.61 86.90 37.41 82.06 94.38 29.81 66.65 84.27 4.45 / 42.20 UGLF Ours 33.93 64.82 74.84 53.10 90.98 97.10 39.66 78.86 91.13 46.59 82.64 96.12 29.59 67.21 85.34 4.51 / 45.06
Unlike existing datasets, TTA videos are collected using handheld cameras by professional analysts from Paralympics Australia. The dataset includes elite level competitions such as the Paralympics and World Para Future. This real-world, single camera setting introduces challenges such as frequent occlusions, motion blur, and dynamic angles—conditions closer to practical applications—unlike the broadcast quality videos used in most prior work. As shown in Figure 6, TTA exhibits the highest event density among all datasets, averaging 7.73 events per 100 frame window, compared to 3.81 in Tennis and significantly fewer in others. This emphasizes the need for models that can capture fine-grained temporal structure and distinguish between visually similar frames within narrow time intervals.
@@ 4.2 Evaluation
Since our proposed module, MSAGSM, is designed to be a lightweight and drop in replacement for GSM, we follow the same training procedure as the E2E-Spot baseline [12], replacing GSM with MSAGSM in the backbone. The model performance is evaluated using the mean Average Precision (mAP), computed by averaging the Average Precision across all event classes. We report mAP under $\delta\in\{0,1,2\}$, where $\delta=0$ requires an exact-frame match and $\delta=1/2$ allow a $\pm 1/\pm 2$-frame tolerance.
@@ 4.3 Implementation Details
We closely follow the training protocol of E2E-Spot [12], adopting its default configurations for a fair comparison. Specifically, all models are trained with input clips of length $L=100$ frames, a batch size of 8, and 5,000 randomly sampled clips per epoch. Training runs for 50 epochs using the AdamW optimizer [21] with a base learning rate of $1\times 10^{-3}$, 3 warm-up epochs, and a cosine decay schedule. As in E2E-Spot, positive classes in the cross-entropy loss are weighted by $w=5$ to handle class imbalance. We also adopt the same backbones—RegNetY-200 [25] and ResNet-18 [10]—from the original framework. Unless otherwise noted, all experiments are conducted on a single NVIDIA L40S GPU.
@@ 4.4 Comparison with State-of-the-Art
To evaluate MSAGSM, we compare against five representative PES models. ASTRA [43] and ASFormer [50] serve as Transformer baselines, while E2E-Spot [12], T-DEED [44], and UGLF [37] are CNN-based methods combining a 2D backbone, temporal module, and decoder. E2E-Spot and UGLF both employ GSM [31], whereas T-DEED adopts GSF [32].
For CNN-based methods, we test both the original temporal module and a drop-in replacement with our MSAGSM, and we also compare different temporal modules within the same backbone–decoder setup. All models are trained from official codebases under identical conditions.
@@
Main Results. Table 2 presents results on five fine-grained sports video datasets under three temporal tolerance thresholds ($\delta=0$, $\delta=1$, $\delta=2$). Across the board, our proposed MSAGSM module consistently outperforms existing temporal modeling approaches. For example, within the E2E-Spot architecture using RegNetY-200 [25], MSAGSM improves upon GSM by +4.09 mAP on Tennis at $\delta=0$, by +3.08 mAP on TTA at $\delta=1$, and by +3.14 mAP on FineDiving at $\delta=1$. Similar performance gains are observed when MSAGSM replaces the original temporal module in more complex models such as T-DEED and UGLF.
@@
Out of all possible comparisons across model backbone configurations, MSAGSM achieves the best result in 37 out of 45 cases when compared directly against other temporal modules. Furthermore, it ranks first in 13 out of 15 columns when considering overall best scores, setting new SOTA performance simply by replacing the temporal component.
@@
Backbone and Module Versatility. To evaluate the generality of our approach, we further experiment with alternative backbones, specifically ResNet-18 [10] in E2E-Spot, T-DEED, and UGLF architectures. For simplicity, the main table includes RegNetY-200 results only, while additional results with ResNet-18 [10] are provided in the Supp. §3.1. For both E2E-Spot and UGLF, which use relatively simple temporal modules, our MSAGSM yields substantial gains with minimal complexity overhead, achieving the best performance in 14/15 and 13/15 evaluation groups, respectively. The improvement is less pronounced for T-DEED, which already employs a more sophisticated temporal design. In this case, adding MSAGSM may introduce redundant or overlapping temporal cues, making feature aggregation less effective.
Efficiency. MSAGSM remains lightweight, requiring only 4.5M parameters and 45.06 GFLOPs in its default setting. This is comparable to existing temporal modules and only slightly larger in size, yet it consistently delivers higher accuracy. Its compact design makes MSAGSM a practical drop-in replacement for both lightweight and large-scale models built on 2D pretrained CNN backbones.
Summary. These results demonstrate the effectiveness of MSAGSM as a robust and efficient temporal modeling module. Its consistent improvements across diverse models and baseline temporal modules, particularly on our proposed TTA benchmark, underscore its suitability for capturing the high-frequency, fine-grained temporal dynamics inherent in sports video understanding.
Table 3: Ablation study on attention head count, temporal distance range, and clip length using the MSAGSM module on the TTA dataset. All results are reported in mAP (%), default clip length=100, best result is in bold.
Setting Value $\delta=1$ $\delta=2$ P(M) / GFLOPs
MSAGSM w/o Attention
E2E-GSM - 66.42 76.90 4.45 / 42.18
Distance [1,2] 66.08 77.31 4.47 / 43.26
Distance [1,2,3] 65.82 74.67 4.49 / 44.32
Distance [1,2,3,4] 64.39 74.91 4.51 / 45.40
GSM with Channel Group Spatial Attention
E2E-GSM - 66.42 76.90 4.45 / 42.18
Heads 1 67.02 75.74 4.46 / 42.56
Heads 2 65.39 72.35 4.47 / 42.92
Heads 3 64.01 71.94 4.47 / 43.28
Varying Number of Attention Heads (Distance = [1,2,3])
Heads 1 60.99 75.38 4.50 / 44.68
Heads 2 69.50 78.12 4.51 / 45.06
Heads 3 65.87 75.87 4.51 / 45.42
Heads 4 61.01 72.98 4.52 / 45.78
@@
Distance [1,2] 60.78 71.10 4.48 / 43.98
Distance [1,2,3] 69.50 78.12 4.51 / 45.06
Distance [1–4] 64.18 71.59 4.53 / 46.12
Distance [1–5] 67.54 67.41 4.55 / 47.18
Distance [1–6] 64.03 75.38 4.57 / 48.26
Varying Clip Length with Distance=[1,2,3], Heads=2
Length 25 56.34 64.85 4.47 / 10.72
Length 50 58.56 66.80 4.47 / 21.46
Length 100 69.50 78.12 4.47 / 47.52
Length 150 62.24 74.50 4.47 / 64.22
Length 200 57.29 66.78 4.47 / 85.50
@@ 4.5 Ablation Study
We conduct ablation studies on the TTA and FineDiving datasets, which represent two contrasting types of sports—one with high event density and the other with sparse events.
Isolating Components. We evaluate the effectiveness of our proposed extensions to GSM by isolating them individually. First, we test the multi-scale GSM with different temporal distance ranges. As shown in Table 3, increasing the range alone leads to inconsistent performance, since exchanging more temporal information without guidance can introduce noise. Next, we assess the channel-group spatial attention GSM by applying different numbers of attention heads to the baseline E2E-GSM. A single head provides a slight gain by directing focus to the most salient region, but additional heads do not yield consistent improvements. This suggests that attention alone lacks sufficient temporal context to fully exploit multiple regions of interest.
Finally, when combining both components, their individual limitations are mitigated: channel-group spatial attention highlights the most informative regions, while multi-scale shifts propagate this focused information across time. Together, they form MSAGSM, which consistently outperforms the baseline.
Attention Heads and Temporal Window Size. We study the effect of varying attention heads (with distance fixed at $[1,2,3]$) and dilation ranges (with heads fixed at two). On TTA, increasing the number of heads from one to two substantially improves mAP (+8.51 and +2.74 at $\delta=1,2$), but adding more heads leads to inconsistent or reduced performance, likely due to over-fragmentation.
For distance ranges, the default $[1,2,3]$ provides the strongest and most stable results. Slight gains are seen with $[1,2,3,4]$, but very large windows (e.g., $1,\dots,6]$) dilute fine-grained cues and hurt stability. Overall, while optimal hyperparameters may vary slightly with dataset characteristics, two attention heads and a temporal window of $[1,2,3]$ yield the most consistent performance across benchmarks on the validation sets. All results in the main tables are reported using these default settings, with further ablations including validations are provided in the Supp. §3.3.
Clip length. We also examine the impact of clip length. As shown in Table 3, a length of 100 frames achieves the best performance, which is consistent with prior findings in [12].
5 Limitations and Future Work